### v2

Abstract
This study discusses the critical issues of Virtual Try-On in contemporary e-commerce and the prospective metaverse, emphasizing the challenges of preserving intricate texture details and distinctive features of the target person and the clothes in various scenarios, such as clothing texture and identity characteristics like tattoos or accessories. In addition to the fidelity of the synthesized images, the efficiency of the synthesis process presents a significant hurdle. Various existing approaches are explored, highlighting the limitations and unresolved aspects, e.g., identity information omission, uncontrollable artifacts, and low synthesis speed. It then proposes a novel diffusion-based solution that addresses garment texture preservation and user identity retention during virtual try-on. The proposed network comprises two primary modules - a warping module aligning clothing with individual features and a try-on module refining the attire and generating missing parts integrated with a mask-aware post-processing technique ensuring the integrity of the individual’s identity. It demonstrates impressive results, surpassing the state-of-the-art in speed by nearly 20 times during inference, with superior fidelity in qualitative assessments. Quantitative evaluations confirm comparable performance with the recent SOTA method on the VITON-HD and Dresscode datasets.
Keywords:
Time efficiency Virtual Try-on Identity retention Virtual Try-on Mask-aware post-processing Diffusion-based networks
Figure 1: Visualization for identity preservation and detail preservation compared with DCV-VTON [6]. DCI-VTON degrades clothes’ texture, struggles with symbols and characters on clothes, and alters tattoos, creating artifacts, while our approach maintains texture and tattoo fidelity.
1 Introduction
Virtual Try-On, which involves placing a garment on a particular individual, holds crucial significance in contemporary e-commerce and the prospective metaverse. The key challenge lies in preserving intricate texture details and distinctive features of the target person, such as appearance and pose. Adapting a garment to different body shapes without altering patterns is particularly challenging, especially when body pose or shape varies significantly.
Recent studies based on deep learning techniques have approached these challenges by defining specific body-garment corresponding regions, particularly addressing obstructions [21] or by adding cloth segmentation information [12]. Another approach is taking advantage of the strength of the Diffusion network, combined with Contrastive Language-Image Pretraining (CLIP) [16], which refines post-warping clothing results along with generating missing body parts [6]. Another method uses implicit warping integrated with Diffusion, guided by CLIP-based networks, to overcome this limitation [23]. All the diffusion-based approaches are proven to surpass the traditional flow-based methods in both quantitative and qualitative assessment [6, 23].
Despite strong generative ability, diffusion-based approaches suffer from extended inference times and uncontrollable artifact generation, affecting the user experience and image fidelity. Meanwhile, another equally important challenge besides the garment texture preservation is retaining the user’s identifying characteristics during virtual try-on - mentioned in [23]. Therefore, to tackle these issues, we propose a novel diffusion approach that not only effectively preserves both the garment texture and identity information but also achieves an impressive inference speed for this task.
Our network comprises two primary modules - a warping module and a try-on module, integrated with post-processing blocks. The warping module is pivotal in aligning clothing with the individual’s features. It considers clothing specifics and person-related information, encompassing key points, dense pose images, and garment type-specific regions of interest (e.g., upper, lower, or full dresses). Subsequently, the try-on module refines the warped attire from the initial module, generating the missing parts of the image. The image then undergoes a conditional post-processing named mask-awareness technique to ensure the fundamental integrity of the individual’s identity. Examples of our impressive results compared to the ones from the SOTA paper [6] are depicted in detail in Fig. 1.
In summary, the main contributions of our work are:
  • We introduce a novel try-on technique to generate photo-realistic results for diverse scenarios.
  • We introduce a novel time-efficient diffusion approach that can adjust and maintain the garment details and generate the missing body parts using sophisticated conditional modules, which effectively guide the model’s focus during the generation process to yield satisfying outcomes.
  • We introduce a mask-aware post-processing technique that not only preserves the individual’s identity details but also improves the overall fidelity of the generated images.
2 Related Work
2.1 Virtual Try-on GAN-based Models
In virtual try-on, achieving high levels of realism and fidelity in garment rendering on digital avatars remains a significant challenge. Traditional deep-learning approaches, primarily utilizing flow-based Generative Adversarial Networks
(GANs), have demonstrated notable potential. Most existing virtual try-on methods follow a two-stage process [12, 4, 21]. The first stage involves a warping module responsible for predicting the appearance flow for the global [12, 4], or local parts [21] of the garment to fit the target person’s pose. This is followed by a second stage, where a GAN-based generator seamlessly integrates the warped garment into the model. Although this method has shown some effectiveness, its heavy reliance on warping quality in the initial stage often leads to less-than-ideal outcomes. This is particularly evident in the realism of garment-skin boundaries and the overall try-on effect in the garment area. Despite their widespread use, these methods have not seen significant innovation to overcome these specific challenges. Therefore, we proposed an approach that aims to push the boundaries of the field by investigating the use of diffusion-based generators.
2.2 Virtual Try-on Diffusion Models
Diffusion models have recently emerged as formidable rivals to GANs in image generation, excelling in producing high-fidelity conditional images. Their application “diffuses” a diverse range of tasks, including text-to-image generation, image-to-image translation, and image editing. Notably, in the context of virtual try-on, diffusion models offer a promising solution by treating the task as a form of image editing conditioned on a specific garment and a full-body image.
A notable diffusion-based approach is presented in TryonDiffusion [23], which leverages a cross-attention mechanism and integrates an implicit warping algorithm with a try-on module. While this method showcases potential in virtual try-on applications, it struggles to retain textural details in the final output.
Other strategies combine the explicit warping module from GAN-based methods with a diffusion model to merge warped clothing and person images. LaDI-VTON [14] and DCI-VTON [6] leverage the capabilities of Stable Diffusion to preserve the texture and details of in-shop garments, achieving high-quality images to a certain level. However, these methods often create unintended artifacts in the final images, which remain challenging to control. Moreover, despite the outstanding generation quality of the diffusion approach, the models’ inference times are excessively long. The major reason lies in the large number of denoising steps (a large number of time steps) during the inference process.
3 Methodology
Addressing these challenges, we propose a novel approach to enhance artifact control. Our method involves developing a diffusion model that not only competes with state-of-the-art(SOTA) models’ performance but also ensures time efficiency. This is achieved by incorporating (un)conditional mask-aware techniques and a modified adding noise algorithm reducing the number of time steps to one, thereby offering a practical and effective solution for virtual try-on applications. The mask-aware techniques are particularly crucial in attaining a realistic wearing effect, as they directly address the common issue of unnatural transitions between clothing and skin, and preserve individual’s identities.
In the virtual try-on task, given an image of a person $I_{p}$ and an image of a garment $I_{g}$, we want to obtain the try-on image that portrays the person wearing the garment. The overall architecture of our approach is depicted in Figure 2. We first pre-process the person image to obtain reference information, including human parsing, 2D pose key points $J_{p}$, and dense pose $I_{dp}$. Following this, our wrapping module utilizes the reference information and produces predicted multi-scale warped-based parsing {$S_{i}$} including both the warped cloth mask $S^{i}_{m}$ and the segmentation of visible body parts $S^{i}_{bp}$, along with multi-scale appearance flows {$f_{i}$}  where the value of $i$ is in range of $[1,7]$ in resolution 512 x 384 and in range of $[1,6]$ in resolution 256 x 192. At the highest resolution level (considered here as $i=1$), these outputs are integrated with the garment and person images, in addition to preserved person information ($M_{p}$), to construct the conditional input ($I_{c}$). The conditional input $I_{c}$ in conjunction with dense pose $I_{dp}$, predicted body part segmentation $S^{i}_{bp}$, garment image $I_{g}$, and person image $I_{p}$ are fed into the try-on module to produce the final image $I_{Out}$.
Figure 2: Overall Generation Pipeline.
3.1 Preprocessing
The procedure begins with generating human parsing maps from the person image ($I_{p}$) by employing advanced human parsing methods [13]. Then, we apply 2D key points [3], and dense pose estimation [7]; their outputs are denoted as $J_{p}$ and $I_{dp}$, respectively. Subsequently, depending on the garment type, we identify and exclude mutable sections based on the human parsing map. In detail, for dress outfits, the changeable parts encompass the entire arm (excluding the hand), the leg, and the exposed body area, including the neck, along with the area unoccupied by the clothing. For upper clothes, this changeable region now combines the arms and torso area of both the person and the clothes. Meanwhile, for lower clothes, the changeable parts are limited to the legs and lower clothing. The remaining segments are combined to create the preserved mask ($M_{p}$).
Figure 3: Warping Module structure. It is crucial to highlight that our model extracts six or seven multi-scale features, depending on the input resolution (i.e., $N$ = 6 or 7). For brevity, the number of scales depicted in this figure is limited to three ($N$ = 3).
3.2 Warping Module
The warping module is the first critical component of our system, designed to handle the task of aligning an in-shop garment image ($I_{g}$) with the corresponding regions on a person image ($I_{p}$). This module is tasked with estimating the warping flow for the garment and identifying the segmentation regions for both the clothing and any occluded body parts. Our warping module adopts almost all the concepts from GP-VTON [21] except internal enhancement. Instead of estimating the local flows for certain parts of the cloth, we target the direct prediction of global flow for warping intact garments likes [12, 9, 8, 5], as depicted in Fig.3 (A). A significant enhancement in our approach is the integration of multi-head cross-attention mechanisms [18] for improved feature integration, which in turn boosts the fidelity of the warping results and overall model efficacy.
$\centering Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d}}V)\ \@add@centering$ (1)
where $Q\in\mathbb{R}^{M\times d},K\in\mathbb{R}^{N\times d},V\in\mathbb{R}^{N\times d}$ represent stacked vectors of query, key, and value, respectively. Here, $M$ is the number of query vectors, $N$ is the number of key and value vectors, and $d$ is the dimension of the vector. In our setup, $Q$ represents the flattened feature of the warped garment $g^{w}_{i}$ the result of warping garment feature $g_{i}$ with input flow $f^{FF}_{in}$, while $K$ and $V$ correspond to the flattened features of the person $p_{i}$. The dot-product-based attention map $\frac{QK^{T}}{\sqrt{d}}V$ serves as an additional feature indicating the similarity between the person and the warped garment. This operation is strategically placed on the Fine Flow Block (FF-B) in the Fusion Block, as depicted in Fig. 3 (B), replacing the concatenation at the feature resolution below 64 x 48 for parameter efficiency. Our warping module detail will be depicted more in the Appendix.
Figure 4: Try-on module training, inference strategy, and details. $S_{bp}$ is the body part segmentation extracted from $I_{p}$.
3.3 Try-on Module
Despite the robust stability exhibited by traditional vanilla diffusion models, they suffer from a notable drawback — prolonged inference times, which can hinder practical applications. However, our variance approach addresses this issue without compromising the model’s generative performance. As indicated in the overview of our strategy in Fig.2, we intend to apply a variant of the modified diffusion model to refine the coarse synthesis result. Following the acquisition of the warping and parsing result, the warped garment is incorporated into the person’s image, alongside the preserved elements and the original background, to form the conditional image $I_{c}$. This image $I_{c}$ serves as the local condition for the try-on diffusion module. Moreover, the global conditions are derived from the target garment image $I_{g}$ using the feature representations extracted by a frozen, pre-trained CLIP [16] image encoder. All are illustrated in Fig. 4(A).
3.3.1 Training pipeline.
During the training process, we work with pair-setting; a noise level is directly added to the ground truth image $I_{gt}$ to create a conditional noise input $I_{n}$. Subsequently, the local condition image $I_{c}$ mentioned earlier is concatenated with the dense pose image $I_{dp}$ dense pose condition, and the conditional noise $I_{n}$. This concatenated input is then fed into the model. The output from the model is then merged with the static components, which include the overlapping areas of the predicted and actual backgrounds as well as the unaltered segment $M_{p}$. This highlights our try-on model’s focus on generating missing image parts and refining clothing details.
The objective function is then applied to the integrated output and the ground truth image. In the optimization process, our model leverages global and local conditions to generate the corresponding inferred person image. In contrast to prior studies related to the diffusion model, where the amount of noise added to the image is adjusted through a time variable, in this study, we use a fixed amount of noise, rendering the time variable redundant during the training process. At this point, our approach could be considered a single-step diffusion model. The adding noise process follows the equation:
$\centering z=z_{0}+\alpha_{n}\epsilon,\@add@centering$ (2)
where $\alpha_{n}$ is the noise ratio and $\epsilon\sim\mathcal{N}(0,I)$. Therefore, our framework is built upon the baseline model of the Denoising Diffusion Implicit Models (DDIM) study [17], with the temporal embedding block supplanted by the CLIP feature.
3.3.2 Inference pipeline.
In the absence of the ground truth image during inference, it becomes imperative to generate an alternative conditional input to facilitate the inference pipeline. In the inference process, the local condition $I_{c}$ is constructed and sent to a post-processing block after the warping module. To be specific, the output of the first post-processing block $I^{\prime}_{c}$ is from combining the local condition image $I_{c}$ with the remaining body part extracted from the ground truth image $I_{p}$. Subsequently, it is made noisy to obtain another noise condition image $I^{\prime}_{n}$. Following a similar diffusion process as outlined in the training phase, the output image $I_{Out}$ is advanced into another post-processing block, culminating in the final result $I^{\prime}_{Out}$. This sequence ensures the generation of a refined output image, compensating for the absence of ground truth data through strategic conditioning and processing stages.
3.3.3 (Un)conditional Post-processing Block.
The post-processing technique, termed (un)conditional mask-aware, operates throughout the entire inference pipeline, ensuring the enhancement of synthesized images with a focus on artifact reduction and detail preservation. The first post-processing technique, unconditional post-processing, aims to enhance the condition image, improving the final synthesized results. Meanwhile, the second technique, conditional post-processing, focuses on minimizing artifacts and preserving identity details.
The unconditional block involves obtaining body segmentation masks $S^{i}_{bp}$ here $i=1$ denoting the largest segmentation map extracted from the warping module and corresponding body masks $S_{bp}$ from the reference person image. Notation that two of these maps have the same resolution. The overlapping mask $S^{\prime}_{bp}$ is used to extract unchanged elements from $I_{p}$, which are then added to $I_{c}$ to get the coarse try-on image $I^{\prime}_{c}$. This image will be then refined by our diffusion model. The conditional post-processing block is used to re-apply the unchanged parts from the input image $I_{p}$ to the refined try-on image $I_{Out}$, to correct any undesirable modification caused by the diffusion model. It has the same structure as the unconditional one, except that it is only applied when the overlapping ratio between $S_{bp}$ and $S^{i}_{bp}$ where $i=1$, is greater than a threshold. Empirically, we found this threshold $R_{overlap}$ needs to be greater than 0.8.
$\centering R_{overlap}=\frac{\left|S^{(1)}_{bp}\bigcap S_{bp}\right|}{S^{(1)}_{bp}}=\frac{S^{\prime}_{bp}}{S^{(1)}_{bp}}\@add@centering$ (3)
Note that we apply this to every single parsing part, including the left/right hand, left/right leg, and the center body part (neck and belly), illustrated in Figure 4 (B) and (C).
3.3.4 Objective function.
In terms of optimization processing for the try-on module, we utilize 2 reconstruction losses which are $\ell_{1}$ loss - $L^{tryon}_{1}$ and perceptual loss - $L^{tryon}_{per}$. Furthermore, an adversarial loss from [11] is also used for better quality results - $L^{tryon}_{adv}$. The total loss for the try-on module can be formulated as:
$\centering L^{tryon}=L^{tryon}_{1}+\alpha^{tryon}_{per}L^{tryon}_{per}+\alpha^{tryon}_{adv}L^{tryon}_{adv}\@add@centering$ (4)
where, $\alpha^{tryon}_{per}$ and $\alpha^{tryon}_{adv}$ are the balance coefficients.
4 Experiments
4.1 Experiments Setting
4.1.1 Dataset.
Our experiments primarily use the VITON-HD dataset [4], comprising 13,679 frontal-view woman and upper clothes image pairs at 1024x768 resolution. Following prior work [4, 12], we split the dataset into a training set of 11,647 pairs and a test set of 2,032 pairs. Experiments are conducted at various resolutions, and we also assess the model on the DressCode dataset [15] for added complexity. Much like VITON-HD [4], the DressCode dataset [15] is a repository of high-quality try-on data pairs, comprising three distinct sub-datasets: dresses, upper-body, and lower-body. In total, the dataset encompasses 53,795 image pairs, distributed across 15,366 pairs for upper-body attire, 8,951 pairs for lower-body clothing, and 29,478 pairs for dresses. To maintain consistency, we apply the same human parsing and key points pose estimation methods used on the DressCode dataset [15] to the VITON-HD dataset [4].
4.1.2 Evaluation Metrics.
In Virtual Tryon evaluations, we consider paired and unpaired settings. Paired assessments use Structural Similarity Index Measure (SSIM) [20] and Learned Perceptual Image Patch Similarity (LPIPS) [22] for image reconstruction, while unpaired settings employ Frechet Inception Distance (FID) [10] and Kernel Inception Distance (KID) [2] to measure the model’s ability to generate new images with changed clothing. Specifically, our evaluation metrics for VITON-HD [4] are all above. Meanwhile, all experiments conducted on the DressCode dataset [15] are executed at a resolution of 512 × 384, and the evaluation metrics only include LPIPS [22], SSIM [20], and FID [10]. In addition, we also measure the speed of synthesizing at the 512 x 384 resolution (T(s)) of the VITON-HD dataset. For a fair comparison in terms of inference time, we ensure consistent configuration settings, employing a single RTX 4090 with a batch size of 4. The model’s inference time is computed by averaging the time taken for end-to-end inference over the entire testing set, repeated 10 times.
4.2 Quantitative Evaluation
In our comparative analysis with existing virtual try-on methods on VTON HD dataset [4] and DressCode dataset [15], including CP-VTON [19], VITON-HD [4], FS-VTON [9], SDAFN [1], PF-AFN [5], HR-VTON [12], GP-VTON [21], and the current state-of-the-art (SOTA) DCI-VTON [6], our method showcases competitive performance across various metrics (Table 1).In terms of VITON HD [4], our model demonstrates compatibility with this SOTA method while DCI-VTON outperforms previous studies in all evaluated categories. Notably, our approach excels in certain aspects while trailing in others, both in paired and unpaired settings. A significant advantage of our model lies in its remarkable inference speed. As detailed in Table 1, our model achieves the best inference times, which is more than 17.43 times faster than DCI-VTON in 512x384 resolution - 1.01s of ours compared to 17.60s of DCI-VTON. Besides, our method also demonstrates superior performance that is compatible with the DCI-VTON across almost all three subsets of DressCode [15].
Table 1: Quantitative comparison with baselines on VITON HD [4] and DressCode [15]. We multiply KID by 100 for better comparison. * is represented for that we only infer the Tryon model. T(s)↓ is the inference times. Bold and underline denote the best and the second best
Method VITON-HD DressCode - (512 x 384)
256 x 192 512 x 384 Upper Lower Dress
LPIPS↓ SSIM↑ FID↓ KID↓ LPIPS↓ SSIM↑ FID↓ KID↓ T(s)↓ LPIPS↓ SSIM↑ FID↓ LPIPS↓ SSIM↑ FID↓ LPIPS↓ SSIM↑ FID↓
CP-VTON [19] 0.089 0.739 30.11 2.034 0.141 0.791 30.25 4.012 - - - - - - - - - -
VITON-HD [4] 0.084 0.811 16.36 0.871 0.076 0.843 11.64 0.3 0.64 - - - - - - - - -
FS-VTON [9] - - - - - - - - - 0.0376 0.9457 13.16 0.0438 0.9381 17.99 0.0745 0.8876 13.87
SDAFN [1] - - - - - - - - - 0.0484 0.9379 12.61 0.0549 0.9317 16.05 0.0852 0.8776 11.8
PF-AFN [5] 0.089 0.863 11.49 0.319 0.082 0.858 11.3 0.283 - 0.038 0.9454 14.32 0.0445 0.9378 18.32 0.0758 0.8869 13.59
HR-VTON [12] 0.062 0.864 9.38 0.153 0.061 0.878 9.9 0.188 1.39 0.0635 0.9252 16.86 0.811 0.9119 22.81 0.1132 0.8642 16.12
GP-VTON [21] - - - - 0.08 0.894 9.2 - - 0.0359 0.9479 11.89 0.042 0.9405 16.07 0.0729 0.8866 12.26
DCI-VTON [6] 0.049 0.906 8.02 0.058 0.043 0.896 8.09 0.028 17.60* 0.0301 - 10.82 0.0348 - 12.41 0.0681 - 12.25
Ours 0.056 0.909 7.53 0.07 0.067 0.909 8.43 0.066 1.01 0.0357 0.9495 10.4 0.0417 0.9413 12.69 0.0727 0.886 11.2
4.3 Ablation Study
By taking 512 x 384 resolution on the VITON-HD dataset as the basic setting, we conduct ablation studies to validate the effectiveness of several components in our networks, and the results are shown in Table 2.
Table 2: Condition effectiveness - ablation study on VITONHD (512x384). We multiply KID by 100 for better comparison
Method VITON-HD – 512 x 384
LPIPS↓ SSIM↑ FID↓ KID↓
w/o Unconditional Postprocessing 0.0706 0.9045 8.52 0.074
w/o Noise Condition Image 0.0923 0.8774 11.00 0.21
w/o Global Condition 0.0759 0.9036 9.11 0.11
w/o Densepose Condition 0.1725 0.8320 24.49 1.43
Ours 0.0675 0.9091 8.43 0.066
Figure 5: Visual condition effectiveness ablation studies for the Tryon module in our approach. Note that all ablation studies are applied to conditional post-processing. Please zoom in for better visualization.
4.3.1 Condition Effectiveness - Tryon Module.
In this ablation study, detailed in Table 2, we systematically assess the components of our Tryon module. First, we explore the impact of removing the unconditional post-processing block. In this experiment, the local condition image $I_{c}$ instead of $I^{\prime}_{c}$ is used to make the noise conditions image, we find it slightly degrades the model quality to imply that unconditional post-processing is still important. Second, we investigate the role of the generated noise condition ($I^{\prime}_{n}$) by substituting it with regular Gaussian noise at the same level. Meanwhile, in the Global Condition effectiveness assessment, we just replaced the CLIP-embedding vector with the same shape zeros vector. In terms of dense pose conditions experiment, we do the same as the global condition case. Our findings underscore the significant influence of both the dense pose condition and the noise condition image, with the former being the most impactful, followed by the latter. Additionally, the global condition - the CLIP-based embedding module shows its importance as the third most influential block. Intriguingly, the removal of the first post-processing block in the Tryon module exhibits minimal impact on model quality. These results are visually demonstrated in Figure 5, where removal of the densepose condition (w/o Densepose Condition) destroys body part structures and clothes, while the exclusion of the noise condition image results in pure noise (w/o Noise Condition Image), undermining clothing texture. The model’s failure to recognize and retain specific cloth textures, evident in the absence of the global condition (w/o Global Condition) - the transparent sleeves or the frill of the cloth, highlights its significance. Interestingly, the decision not to apply unconditional post-processing results in only a slight reduction in output detail (w/o Unconditional Post-processing).
Figure 6: Qualitative comparison with baseline in VITON-HD Dataset [4] at 512 x 384 resolutions. Please zoom in for better visualization.
4.3.2 Conditional Mask-aware Post-processing Performance.
In addition, we not only evaluate the effectiveness of our "plug-and-play" post-processing block for our study but also do on previous studies that have the same network structure to prove its efficiency. Table 3 illustrates that our mask-aware postprocessing block can enhance the model performance in not only our study but also all previous studies under any constraint of resolution. Note that, the result in this table is experienced based on the pre-trained model provided by the research author, which is run on the same setting as in our inference time calculation.
Table 3: Conditional Post-processing on ours and other previous studies. The * represents that we only ran the experiment based on the pre-trained model at 1024 x 768 resolution published by the authors of these studies, then plugged in our post-processing technique. Our approach was applied to images of 512x384 resolution, consistent with the resolution used in the referenced experiments. We multiply KID by 100 for better comparison.
Method VITON-HD
LPIPS↓ SSIM↑ FID↓ KID↓
VITON-HD* + w/o postprocessing 0.1308 0.8665 11.755 0.284
VITON-HD* + postprocessing 0.1296 0.8691 11.754 0.282
HR-VTON* + w/o postprocessing 0.1106 0.8816 11.204 0.268
HR-VTON* + postprocessing 0.1068 0.8877 11.195 0.264
Ours + w/o postprocessing 0.0706 0.9016 8.49 0.071
Ours + postprocessing 0.0675 0.9091 8.43 0.066
4.4 Qualitative Evaluation
In Fig. 6, we showcase composite images from various methods on the VITON-HD Dataset at 512 x 384 resolution. Our approach demonstrates superior performance to DCI-VTON, the current SOTA results on the VITON-HD dataset in terms of detail features. Our architecture generates highly realistic images, surpassing previous studies like VITON-HD and HR-VTON, particularly excelling in costume details. Notably, our model effectively handles challenging cases such as see-through collars - the first example, where DCI-VTON falls short. In intricate scenarios like crossed arms, both DCI-VTON and our model perform well, but our method excels in preserving finer outfit details (the neckline color is white in DCI-VTON and blue in ours). The last row exemplifies DCI-VTON introducing artifacts (the shirt strap in the neck part of the DCI-VTON result), while our approach maintains realism without these artifacts.
Our research not only preserves outfit details and minimizes artifacts but also addresses the crucial issue of retaining identity information. Leveraging our conditional mask-awareness post-processing technique, we successfully address this concern. Figure 7 illustrates the efficacy of our post-processing, comparing results before and after its application alongside the outcomes of the SOTA paper under the same conditions. Our technique effectively preserves identity information, demonstrated in cases where DCI-VTON fails to retain arm tattoos (first and last rows). DCI-VTON introduces artifacts in the neck area (strange necklaces) and the lower-cloth context, compromising image authenticity ($3^{rd}$, $4^{th}$, and $5^{th}$ rows ). Additionally, DCI-VTON struggles to maintain outfit details, as seen in the first line and highlighted regions (purple square) in lines 4 and 5. Our research ensures the maximum retention of identity information for the inference person without compromising output authenticity.
Figure 7: Example for the identity preservation on VTON-HD Dataset (512 x 384). The blue circle represents the important parts that need to be preserved, the green rectangle is for the wrong context or the artifact made by DCI-VTON compared to our approach, and the red rectangle is for the detailed part of the clothes that can not be retained compared to our approach. Please zoom in for better visualization.
5 Conclusions
In this study, we have addressed the challenges of Virtual Try-On technology, by introducing a novel diffusion-based approach that adeptly preserves both garment texture and user identity. Our integrated system, comprising a warping module and a try-on module, enhanced with a mask-awareness post-processing technique, significantly outperforms existing methods in terms of inference speed, being over 17.43 times faster than the current state-of-the-art, while also maintaining superior fidelity in output. This dual focus on efficiency and detail preservation marks a substantial advancement in the field. However, it is important to acknowledge a limitation in our approach: the necessity for an elaborate post-processing process. This additional step, while crucial for ensuring the integrity of the individual’s identity and the garment’s texture, adds complexity to the overall system. Despite this, the proposed method presents a promising solution for real-world applications, offering a more seamless and accurate virtual try-on experience. Future work could aim to streamline this post-processing phase, further enhancing the system’s efficiency and applicability in diverse scenarios, including a wider range of clothing styles and body types.